OpenStudy (anonymous):

find dx/dt when y=squareroot(9-2x^2) given x=2 y=1

OpenStudy (anonymous):

the derivative of \[\sqrt{9-2x^2}\] is \[\frac{1}{2\sqrt{9-2x^2}}\times -4x\] by the chain rule bit of algebra gets \[-\frac{2x}{\sqrt{9-2x^2}}\] now replace \(x\) by \(2\)